OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2

OpenStudy (anonymous):

YOU KNOW HOW TO DO THIS !?

OpenStudy (anonymous):

Use Distributive property. -4(2x-1) = -8x+4 because negative and minus become positive. therefore, u change the sign minus into plus. so you know to solve it.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think the answer is 5

OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2 -8x + 4 = -6x -12 -2 -8x + 4 = -6x - 14 -2x = -18 x = 9

OpenStudy (anonymous):

look. -8x+4 = (-6x-12)-2 So You want to get rid of x. So subtract -6x to -8x, and you will get -2. -2x+4=-14 because -12-2=-14. Then u subtract 4 from -14. You get -2x=-18. Then you divide -2 from -18 and you get x=9. because when you divide both negative numbers, it become positive.